# Relational Conformal Prediction for Correlated Time Series

### v1

@@ Abstract
We address the problem of uncertainty quantification in time series forecasting by exploiting observations at correlated sequences. Relational deep learning methods leveraging graph representations are among the most effective tools for obtaining point estimates from spatiotemporal data and correlated time series. However, the problem of exploiting relational structures to estimate the uncertainty of such predictions has been largely overlooked in the same context. To this end, we propose a novel distribution-free approach based on the conformal prediction framework and quantile regression. Despite the recent applications of conformal prediction to sequential data, existing methods operate independently on each target time series and do not account for relationships among them when constructing the prediction interval. We fill this void by introducing a novel conformal prediction method based on graph deep learning operators. Our method, named CoRel (CoRel), does not require the relational structure (graph) to be known as a prior and can be applied on top of any pre-trained time series predictor. Additionally, CoRel includes an adaptive component to handle non-exchangeable data and changes in the input time series. Our approach provides accurate coverage and archives state-of-the-art uncertainty quantification in relevant benchmarks.
@@ 1 Introduction
Many recent advancements in deep learning methods for time series forecasting rely on learning from large collections of (related) time series [1, 2]. In many application domains, such time series are characterized by a rich spatiotemporal dependency structure that can be exploited by introducing inductive biases in the forecasting architecture [3], to steer the learning procedure toward the most plausible models. Accounting for the existing dependencies, usually represented as a graph, allows the resulting models to obtain accurate predictions with a reduced sample complexity [4, 3]. Besides the accuracy of the point estimates, the reliability of the forecasts is a critical aspect of the problem and a key element to enable effective decision-making in many applications [5, 6]. Uncertainty quantification methods [7, 8] can improve reliability by providing confidence intervals on the forecasting error magnitude, allowing for making more informed decisions [9]. This is particularly true for risk-sensitive applications such as healthcare [10] and load forecasting [11]. In this context, inter-series (spatiotemporal) dynamics offer both a challenge and an opportunity. Indeed, while these dependencies can lead to wide prediction intervals (PIs) if overlooked, they may also provide additional knowledge to reduce uncertainty [12].
@@
Existing probabilistic forecasting frameworks often rely on strong distributional assumptions and major modifications of the base point predictor [1, 13]. As such, they cannot be used to quantify uncertainty given a pre-trained forecasting model. In such a setting, conformal prediction (CP) [8, 14] methods are particularly appealing. CP is an uncertainty quantification framework that estimates confidence intervals with marginal coverage guarantees from observed prediction residuals. One of the main assumptions of standard CP methods is that of exchangeability between the data used to estimate the confidence intervals and the test data points, i.e., the assumption that their joint probability distribution is invariant to the ordering of the associated sequence of random variables [15]. Although this assumption does not usually hold when operating on time series [16], several methods have successfully adapted CP to estimate forecast uncertainty [17, 18, 19, 20, 21]. Nevertheless, existing CP approaches operate on each (possibly multivariate) time series independently and cannot account for dependencies among correlated time series.
@@
In this paper, we propose CoRel, a novel CP approach leveraging graph representations and graph deep learning (GDL) for quantifying uncertainty in correlated time series forecasting. In our framework, a spatiotemporal graph neural network (STGNN) [4, 3] is trained on a calibration set to approximate the quantile function of the distribution of prediction residuals. Relationships among time series, assumed to be sparse, are learned end-to-end from the observed residuals owing to a graph structure learning module integrated into the processing. Our approach estimates the error quantile function for each time series at each time step, by conditioning the shared uncertainty quantification model on past observations at neighboring nodes (as defined by the learned graph structure). Finally, an adaptive component is added to handle potential non-stationarities by relying on a small set of parameters specific to each time series. Our approach can be applied to the residuals generated by any point forecasting model, even those that completely disregard potential relationships among the input time series.
@@
  • A family of graph-based architectures to estimate uncertainty that share most of the learnable parameters among the processed time series, while including node-level parameters that dynamically adapt to changes in each target sequence.
@@ 2.1 Problem formulation
Consider a collection of $N$ sparsely correlated time series. Denote by ${\bm{x}}^{i}_{t}\in{\mathbb{R}}$ the scalar target variable associated with the $i$-th time series at time step $t$; ${\bm{X}}_{t}\in{\mathbb{R}}^{N\times 1}$ indicates the $N$ stacked target variables w.r.t. the entire time series collection. ${\bm{X}}_{t:t+T}$ indicates the sequence within time interval $[t,t+T)$; conversely, with the shorthand ${\bm{X}}_{<t}$ refers to observations up to time step $t$ (excluded). Time series are assumed to be homogenous, i.e., all the variables (observables) describe the same physical quantity (e.g., temperature or energy consumption). Analogously, ${\bm{U}}_{t}\in{\mathbb{R}}^{N\times d_{u}}$ indicates the $d_{u}$-dimensional exogenous covariates associated with each time series. We assume that the $i$-th time series is generated by a stochastic time-invariant process such as
@@
Let us hypothesize the existence of a sparse predictive causality à la Granger [22], i.e., we assume that the values of a single time series are related to the values of a (small) subset of other time series in the collection. The extension of the framework to collections of multivariate time series is straightforward, but we focus on the univariate case to simplify the presentations. The problem of dealing with non-stationary processes will be discussed in Sec. 3.4.
@@
We are interested in a model that makes point forecasts by predicting the unknown $H$-steps-ahead ($H\geq 0$) observation ${\bm{X}}_{t+H}$ given a window $W\geq 1$ of past observations ${\bm{X}}_{t-W:t}$ and the associated exogenous variables ${\bm{U}}_{t-W:t}$ as
@@
$P\left({\bm{x}}^{i}_{t}\in C^{\alpha}_{i,t}\left(\widehat{\bm{X}}_{t+H}\right)\right)\geq 1-\alpha,$ (3)
where $\alpha$ is the desired confidence level. If the interval satisfies Eq. 3, we say that the PI achieves marginal coverage $1-\alpha$. Similarly, we say that the PI provides conditional coverage if
$P\left({\bm{x}}^{i}_{t}\in C^{\alpha}_{i,t}\left(\widehat{\bm{X}}_{t+H}\right)\Big{|}{\bm{X}}_{<t},{\bm{U}}_{<t}\right)\geq 1-\alpha.$ (4)
Conditional coverage provides stronger guarantees and it is often harder to achieve [15]. In the following, we will omit the dependence of the interval on the forecasts and simply write $C^{\alpha}_{i,t}$. Among uncertainty quantification methods, we are interested in post-hoc approaches that can build confidence intervals for any given pre-trained point predictor ${\mathcal{F}}_{\theta}$ without requiring any modification of the base forecasting architecture.
@@ 2.2 Conformal prediction
As anticipated in Sec. 1, standard CP methods [8, 14] are a class of distribution-free uncertainty quantification techniques that build PIs from empirical quantiles of conformal scores. In the forecasting setting, we consider as conformal scores the prediction residuals,
@@
and use ${\bm{R}}_{t}$ to denote residuals w.r.t. the entire time series collection. Under appropriate assumptions, CP methods can build valid and informative PIs [14, 16]. SPLIT CONFORMAL PREDICTION (SCP) [8] is arguably the most common approach and exploits scores computed on a calibration set that is disjoint from the training data (i.e., a post-hoc approach).
As mentioned, most standard CP methods rely on the assumption that calibration and test data are exchangeable which allows the procedure to treat them symmetrically and obtain valid (marginal) coverage guarantees [15]. Since this assumption does not hold when dealing with time series data, there have been several recent results extending the CP framework beyond exchangeability [23, 17, 24, 18]. In particular, Barber et al. [16] showed that approximate coverage can be achieved by reweighting the residuals to account for the lack of exchangeability between calibration and test set. Auer et al. [21] learn such a reweighting scheme through an attention-based architecture. Differently, Xu and Xie [19] introduce SCPI, a method based on fitting a quantile random forest [25] on the most recent prediction residuals at each time step. Similar to SCPI, our approach relies on quantile regression to build PIs but differently from existing methods, it exploits observations in arbitrary sets of time series by relying on GDL operators.
@@ 2.3 Quantile regression
Quantile regression [26] is an established statistical framework that consists of learning a model of the quantile function (the inverse c.d.f.) of a target distribution from observations. In particular, given ${\bm{y}}\sim p({\bm{y}}|{\bm{x}})$ and observations $(x_{1},y_{1}),\dots,(x_{N},y_{N})$, a standard approach to estimate the $\alpha$-quantile is to train a model by minimizing the so-called pinball loss
@@
Quantile regression has been incorporated in several probabilistic forecasting architectures [1]. The simplest approach consists of using a multi-output network to predict a set of quantiles of interest and interpolate among them to approximate the entire quantile function [27]. More complex approaches rely on, e.g., splines [28]. Conversely, implicit quantile networks (IQNs) [29, 30, 31] approximate the quantile function by being trained to minimize the loss in Eq. 6 given the quantile level $\alpha$ as input and sampling a random $\alpha$ for each sample in a mini-batch.
@@ 2.4 Graph deep learning for time series forecasting
Graph neural networks (GNNs) [32, 33] process graph-structured data by incorporating the graph topology as an inductive bias, e.g., by relying on message-passing layers [34]. STGNNs [4, 3] leverage message-passing layers within sequence modeling architectures to process spatiotemporal data and collections of time series where dependencies are represented as a (possibly dynamic) graph. We consider as reference architectures time-then-space (TTS) models [35, 3] where each time series in the collection is processed independently from the others by a temporal encoder whose output is then fed into a stack of GNN layers. In particular, we adopt the following template architecture:
@@
$\hat{\bm{y}}_{t}$ $=\textsc{Readout}\left({\bm{h}}_{t}^{L}\right),$ (9)
where ${\bm{A}}\in{\mathbb{R}}^{N\times N}$ is the graph adjacency matrix and $\hat{\bm{y}}_{t}$ a generic node-level prediction associated with the problem at hand. $\textsc{SeqEnc}({}\cdot{})$ and $\text{GNN}_{l}({}\cdot{})$ denote, respectively, any sequence modeling architecture, e.g., an RNN, and any GNN layer, e.g., based on message-passing. Representations can then be mapped into predictions $\widehat{\bm{Y}}_{t}$ by using any readout block, e.g., a multilayer perceptron (MLP). STGNNs have been used as forecasting architecture (${\bm{Y}}_{t}={\bm{X}}_{t+H}$) with great success. In the following, we will exploit this framework as a backbone for estimating the residual quantile distribution. We refer to Sec. 4 and Jin et al. [4] for more discussion on the application of STGNNs in the context of time series analysis.
@@ 3 Conformal relational prediction
Our objective is to build PIs by exploiting relational dependencies across the residuals of the target time series. We model the dependencies as edges of a graph and learn them under the assumption that the relational structure is sparse, which reduces the computational costs and act as an inductive bias on the structure learning architecture. By relying on such representation, we can leverage GDL methods for time series to process the data. In particular, we train a STGNN on the residuals of the calibration set to predict the quantiles of the error distribution. Conditioning the prediction on the recent history of related time series allows for taking the dependency structure of the data into account when estimating uncertainty: a key aspect in applying conformal prediction to non-exchangeable data [16]. Compared to existing methods [19, 21] that only capture temporal dependencies, our approach allows for modeling spatiotemporal dependencies among different time series. Sec. 3.1 presents the details of the proposed conformal inference procedure by assuming that the relational structure at each time step is defined by an adjacency matrix ${\bm{A}}\in{\mathbb{R}}^{N\times N}$ (Sec. 3.1). We then show how to learn the graph structure directly from data and make the model adaptive in Sec. 3.2 and Sec. 3.4, respectively. Finally, we discuss the theoretical properties of the approach in Sec. 3.3. Fig. 1 shows an overview of the architecture.
@@ 3.1 Relational Quantile Predictor
Consider a standard SCP setup, where the training data are split into training and calibration sets. For the moment, we disregard possible nonstationarities in the data considering the problem setup introduced in Sec. 2.1 and encode spatial dependencies in the adjacency matrix ${\bm{A}}\in{\mathbb{R}}^{N\times N}$. While the training set is used to fit the point predictor ${\mathcal{F}}_{\theta}$, we use the prediction residuals in the calibration set (${\mathcal{R}}^{cal}$) to learn the quantile function of the error distribution at each time step.
We implement the quantile regressor as a hybrid global-local STGNN, which mixes global (shared) parameters with local, target-specific components [36]. Sharing most learnable parameters across all time series reduces sample complexity, while local parameters allow for processing each series differently. Specifically, we keep all processing blocks shared and associate a learnable node embedding ${\bm{v}}^{i}\in{\mathbb{R}}^{d_{v}}$ with each time series [37]. More specifically, our model is a quantile network (see Sec. 2.3) composed of the following processing layers:
@@
$\hat{\bm{q}}^{\alpha,i}_{t+H}$ $=\textsc{QDec}\left(\alpha,{\bm{z}}_{t}^{i},{\bm{v}}^{i}\right),$ (12)
where ${\bm{r}}^{i}_{t-1}$ are prediction residuals (Eq. 5) and $\hat{\bm{q}}^{\alpha,i}_{t+H}$ is the predicted $\alpha$-quantile at time step $t+H$ for the $i$-th time series. $\textsc{Enc}({}\cdot{})$ denotes any encoding layer,e.g., a linear transformation or an MLP. For the STGNN block, several designs are possible (e.g., see Jin et al. [4]); the one we follow is the template in Sec. 2.4. $\textsc{QDec}({}\cdot{})$ is a readout mapping the representations at each node to the prediction of the quantile of specified level $\alpha$. We refer to the family of quantile networks defined in Eq. 1012 as relational quantile predictors (RelQPs) and use the notation
@@
where ${\mathcal{Q}}_{\psi}$ indicates the shared (global) part of the network and $\widehat{\bm{Q}}^{\alpha}_{t}\in{\mathbb{R}}^{N}$ denotes the predicted $\alpha$-quantiles at time step $t$ w.r.t. the full time series collection. Note that the framework can easily accommodate further inputs at the encoding block (e.g., we can condition the regression on ${\bm{X}}_{<t}$ and ${\bm{U}}_{<t}$). The model is trained by, at each time step in the calibration set, the pinball loss (Eq. 6) w.r.t. the full-time series collection. Through the message-passing layers, the residuals of each time series contribute to estimating the quantiles of the error distribution at neighboring nodes. In practice, we restrict the input of the regressor to the most recent observations rather than considering the full sequence (the window length here can also be different from the one used by the point predictor).
@@
$\widehat{C}^{\alpha}_{t}=\left[\widehat{\bm{X}}_{t+H}+\widehat{{\bm{Q}}}^{\alpha/2}_{t},\widehat{\bm{X}}_{t+H}+\widehat{{\bm{Q}}}^{1-\alpha/2}_{t}\right],$ (14)
@@
$\hat{\beta}$ $=\operatorname*{arg\,min}_{\beta}\Big{|}\widehat{{\bm{Q}}}^{1-\alpha/2+\beta}_{t}-\widehat{{\bm{Q}}}^{\alpha/2+\beta}_{t}\big{|}$ (15)
$\widehat{C}^{\alpha}_{t}$ $=\left[\widehat{\bm{X}}_{t+H}+\widehat{{\bm{Q}}}^{\alpha/2+\hat{\beta}}_{t},\widehat{\bm{X}}_{t+H}+\widehat{{\bm{Q}}}^{1-\alpha/2+\hat{\beta}}_{t}\right],$ (16)
where $\widehat{C}^{\alpha}_{t}$ indicates the estimated PI. While both Eq. 14 and Eq. 16 correspond to the same confidence level, the PI in Eq. 16 can be narrower, at the expense of the additional computation needed to obtain $\hat{\beta}$ [18]. In practice, one can choose between the two approaches given computational constraints and the difference in performance observed on a validation set. Note that $\widehat{C}^{\alpha}_{i,t}$ can provide only approximate coverage, as it is subject to approximation errors of the true quantile function. Sec. 3.3 will discuss this aspect in detail. A potential drawback of CoRel is that residuals cannot be assumed exchangeable in most practical scenarios. The error distribution can be non-stationary, making it difficult to obtain any coverage guarantee. To mitigate the problem, Sec. 3.4 discusses an efficient and scalable approach to make the framework adaptive by updating local components of the architecture over time. Finally, it is worth noting that the use of relational components in CoRel relies on the actual presence of the associated dependencies in the data. In practical applications, the presence of spatiotemporal correlations in the residuals can be verified through ad-hoc statistical tests  [12, 38], whose outcome can support the adoption of CoRel.

### v2

@@ Abstract
We address the problem of uncertainty quantification in time series forecasting by exploiting observations at correlated sequences. Relational deep learning methods leveraging graph representations are among the most effective tools for obtaining point estimates from spatiotemporal data and correlated time series. However, the problem of exploiting relational structures to estimate the uncertainty of such predictions has been largely overlooked in the same context. To this end, we propose a novel distribution-free approach based on the conformal prediction framework and quantile regression. Despite the recent applications of conformal prediction to sequential data, existing methods operate independently on each target time series and do not account for relationships among them when constructing the prediction interval. We fill this void by introducing a novel conformal prediction method based on graph deep learning operators. Our approach, named CoRel (CoRel), does not require the relational structure (graph) to be known a priori and can be applied on top of any pre-trained predictor. Additionally, CoRel includes an adaptive component to handle non-exchangeable data and changes in the input time series. Our approach provides accurate coverage and achieves state-of-the-art uncertainty quantification in relevant benchmarks.
uncertainty quantification, time series forecasting, conformal prediction, graph deep learning, correlated time series
@@ 1 Introduction
Many recent advancements in deep learning methods for time series forecasting rely on learning from large collections of (related) time series (Benidis et al., 2022; Liang et al., 2024). In many application domains, such time series are characterized by a rich spatiotemporal dependency structure that can be exploited by introducing inductive biases in the forecasting architecture (Cini et al., 2023a), to steer the learning procedure toward the most plausible models. Accounting for the existing dependencies, usually represented as a graph, allows the resulting models to obtain accurate predictions with a reduced sample complexity (Jin et al., 2023; Cini et al., 2023a). Besides the accuracy of the point estimates, the reliability of the forecasts is a critical aspect of the problem and a key element to enable effective decision-making in many applications (Makridakis, 1996; Petropoulos et al., 2022). Uncertainty quantification methods (Smith, 2024; Vovk et al., 2005) can improve reliability by providing confidence intervals on the forecasting error magnitude, allowing for making more informed decisions (Hyndman and Athanasopoulos, 2018). This is particularly true for risk-sensitive applications such as healthcare (Makridakis et al., 2019) and load forecasting (Gasparin et al., 2022). In this context, inter-series (spatiotemporal) dynamics offer both a challenge and an opportunity. Indeed, while these dependencies can lead to wide prediction intervals (PIs) if overlooked, they may also provide additional knowledge to reduce uncertainty (Zambon and Alippi, 2022).
@@
Existing probabilistic forecasting frameworks often rely on strong distributional assumptions and major modifications of the base point predictor (Benidis et al., 2022; Salinas et al., 2020). As such, they cannot be used to quantify uncertainty given a pre-trained forecasting model. In such a setting, conformal prediction (CP) (Vovk et al., 2005; Angelopoulos et al., 2023) methods are particularly appealing. CP is an uncertainty quantification framework that estimates confidence intervals with marginal coverage guarantees from observed prediction residuals. One of the main assumptions of standard CP methods is that of exchangeability between the data used to estimate the confidence intervals and the test data points, i.e., the assumption that their joint probability distribution is invariant to the ordering of the associated sequence of random variables (Angelopoulos et al., 2024). Although this assumption does not usually hold when operating on time series (Barber et al., 2023), several methods have successfully adapted CP to estimate forecast uncertainty (Stankeviciute et al., 2021; Xu and Xie, 2023a, b; Jensen et al., 2022; Auer et al., 2023). Nevertheless, existing CP approaches operate on each (possibly multivariate) time series independently and cannot account for dependencies among correlated time series.
@@
In this paper, we propose CoRel, a novel CP approach leveraging graph representations and graph deep learning (GDL) for quantifying uncertainty in correlated time series forecasting. In our framework, a spatiotemporal graph neural network (STGNN) (Jin et al., 2023; Cini et al., 2023a) is trained on a calibration set to approximate the quantile function of the distribution of prediction residuals. Relationships among time series, assumed to be sparse, are learned end-to-end from the observed residuals owing to a graph structure learning module integrated into the processing. Our approach estimates the error quantile function for each time series at each time step, by conditioning the shared uncertainty quantification model on past observations at neighboring nodes (as defined by the learned graph structure). Finally, an adaptive component is added to handle potential non-stationarities by relying on a small set of parameters specific to each time series. Our approach can be applied to the residuals generated by any point forecasting model, even those that completely disregard potential relationships among the input time series.
@@
  • A family of graph-based architectures to estimate uncertainty that shares most of the learnable parameters among the processed time series, while including node-level parameters that dynamically adapt to changes in each target sequence.
@@ 2 Preliminaries
This section introduces the problem settings and the preliminary concepts that serve as foundations for our approach.
@@ 2.1 Problem Formulation
Consider a collection of $N$ correlated time series. Denote by ${\bm{x}}^{i}_{t}\in{\mathbb{R}}$ the scalar target variable associated with the $i$-th time series at time step $t$; ${\bm{X}}_{t}\in{\mathbb{R}}^{N\times 1}$ indicates the $N$ stacked target variables w.r.t. the entire time series collection. ${\bm{X}}_{t:t+T}$ indicates the sequence within time interval $[t,t+T)$; conversely, with the shorthand ${\bm{X}}_{<t}$ refers to observations up to time step $t$ (excluded). Time series are assumed to be homogenous, i.e., all the variables (observables) describe the same physical quantity (e.g., temperature or energy consumption). Analogously, ${\bm{U}}_{t}\in{\mathbb{R}}^{N\times d_{u}}$ indicates the $d_{u}$-dimensional exogenous covariates associated with each time series. We assume that the $i$-th time series is generated by a stochastic time-invariant process such as
@@
Let us hypothesize the existence of a sparse predictive causality à la Granger (Granger, 1969), i.e., we assume that the values of a single time series are related to the values of a (small) subset of other time series in the collection. The extension of the framework to collections of multivariate time series is orthogonal to the proposed approach (e.g., see Feldman et al. 2023); we focus on the univariate case to maintain a contained scope. The problem of dealing with non-stationary processes will be discussed in Sec. 3.4.
@@
We are interested in a model that produces point forecasts by predicting the unknown $H$-steps-ahead ($H\geq 0$) observation ${\bm{X}}_{t+H}$ given a window $W\geq 1$ of past observations ${\bm{X}}_{t-W:t}$ and the associated exogenous variables ${\bm{U}}_{t-W:t}$ as
@@
$P\left({\bm{x}}^{i}_{t+H}\in C^{\alpha}_{i,t}\left(\widehat{\bm{X}}_{t+H}\right)\right)\geq 1-\alpha,$ (3)
where $\alpha$ is the desired confidence level. If the interval satisfies Eq. 3, we say that the PI achieves marginal coverage $1-\alpha$. Similarly, we say that the PI provides conditional coverage $1-\alpha$ if
$P\left({\bm{x}}^{i}_{t+H}\in C^{\alpha}_{i,t}\left(\widehat{\bm{X}}_{t+H}\right)\Big{|}{\bm{X}}_{<t},{\bm{U}}_{<t}\right)\geq 1-\alpha.$ (4)
Conditional coverage provides stronger guarantees and it is often harder to achieve (Angelopoulos et al., 2024). In the following, we will omit the dependence of the interval on the forecasts and simply write $C^{\alpha}_{i,t}$. Among uncertainty quantification methods, we are interested in post-hoc approaches that can build confidence intervals for any given pre-trained point predictor ${\mathcal{F}}_{\theta}$ without requiring any modification of the base forecasting architecture.
@@ 2.2 Conformal Prediction
As anticipated in Sec. 1, standard CP methods (Vovk et al., 2005; Angelopoulos et al., 2023) are a class of distribution-free uncertainty quantification techniques that build PIs from empirical quantiles of conformal scores. In the forecasting setting, we consider as conformal scores the prediction residuals,
@@
and use ${\bm{R}}_{t}$ to denote residuals w.r.t. the entire time series collection. Under appropriate assumptions, CP methods can build valid and informative PIs (Angelopoulos et al., 2023; Barber et al., 2023). SPLIT CONFORMAL PREDICTION (SCP) (Vovk et al., 2005) is arguably the most common approach and exploits scores computed on a calibration set that is disjoint from the training data (i.e., a post-hoc approach).
As mentioned, most standard CP methods rely on the assumption that calibration and test data are exchangeable, which allows the procedure to treat them symmetrically and obtain valid (marginal) coverage guarantees (Angelopoulos et al., 2024). Since this assumption does not hold when dealing with time series data, there have been several recent results extending the CP framework beyond exchangeability (Tibshirani et al., 2019; Stankeviciute et al., 2021; Gibbs and Candes, 2021; Xu and Xie, 2023a). In particular, Barber et al. (2023) showed that approximate coverage can be achieved by reweighting the residuals to account for the lack of exchangeability between calibration and test set. Auer et al. (2023) learn such a reweighting scheme through an attention-based architecture. Differently, Xu and Xie (2023b) introduce SPCI, a method based on fitting a quantile random forest (Meinshausen and Ridgeway, 2006) on the most recent prediction residuals at each time step. Similar to SPCI, our approach relies on quantile regression to build PIs but differently from existing methods, it exploits observations in arbitrary sets of time series by relying on GDL operators.
@@ 2.3 Quantile Regression
Quantile regression (Koenker and Hallock, 2001) is an established statistical framework that consists of learning a model of the quantile function (the inverse c.d.f.) of a target distribution from observations. In particular, given ${\bm{y}}\sim p({\bm{y}}|{\bm{x}})$ and observations $(x_{1},y_{1}),\dots,(x_{N},y_{N})$, a standard approach to estimate the $\alpha$-quantile is to train a model by minimizing the so-called pinball loss
@@
Quantile regression has been incorporated in several probabilistic forecasting architectures (Benidis et al., 2022). The simplest approach consists of using a multi-output network to predict a set of quantiles of interest and interpolate among them to approximate the entire quantile function (Wen et al., 2017). More complex approaches rely on, e.g., splines (Gasthaus et al., 2019). Conversely, implicit quantile networks (IQNs) (Dabney et al., 2018; Ostrovski et al., 2018; Gouttes et al., 2021) approximate the quantile function by being trained to minimize the loss in Eq. 6 given the quantile level $\alpha$ as input and sampling a random $\alpha$ for each sample in a mini-batch.
@@ 2.4 Graph Deep Learning for Time Series Forecasting
Graph neural networks (GNNs) (Bacciu et al., 2020; Bronstein et al., 2021) process graph-structured data by incorporating the graph topology as an inductive bias, e.g., by relying on message-passing layers (Gilmer et al., 2017). STGNNs (Jin et al., 2023; Cini et al., 2023a) leverage message-passing layers within sequence modeling architectures to process spatiotemporal data and collections of time series where dependencies are represented as a (possibly dynamic) graph. We consider as reference architectures time-then-space (TTS) models (Gao and Ribeiro, 2022; Cini et al., 2023a) where each time series in the collection is processed independently from the others by a temporal encoder whose output is then fed into a stack of GNN layers. In particular, we adopt the following template architecture:
@@
$\hat{\bm{y}}^{i}_{t}$ $=\textsc{Readout}\left({\bm{h}}_{t}^{i,L}\right),$ (9)
where ${\bm{A}}\in{\mathbb{R}}^{N\times N}$ is the graph adjacency matrix and $\hat{\bm{y}}^{i}_{t}$ a generic node-level prediction associated with the problem at hand. $\textsc{SeqEnc}({}\cdot{})$ and $\text{GNN}_{l}({}\cdot{})$ denote, respectively, any sequence modeling architecture, e.g., an RNN, and any GNN layer, e.g., based on message-passing. Representations can then be mapped into predictions $\widehat{\bm{Y}}_{t}$ by using any readout block, e.g., a multilayer perceptron (MLP). STGNNs have been used as forecasting architecture (${\bm{Y}}_{t}={\bm{X}}_{t+H}$) with great success. In the following, we will exploit this framework as a backbone for estimating the residual quantile distribution. We refer to Sec. 4 and Jin et al. (2023) for more discussion on the application of STGNNs in the context of time series analysis.
@@ 3 Conformal Relational Prediction
Our objective is to build PIs by exploiting relational dependencies across the residuals of the target time series. We model the dependencies as edges of a graph and learn them under the assumption that the relational structure is sparse, which reduces the computational costs and act as an inductive bias on the structure learning architecture. By relying on such representation, we can leverage GDL methods for time series to process the data. In particular, we train a STGNN on the residuals of the calibration set to predict the quantiles of the error distribution. Conditioning the prediction on the recent history of related time series allows for taking the dependency structure of the data into account when estimating uncertainty: a key aspect in applying conformal prediction to non-exchangeable data (Barber et al., 2023). Compared to existing methods (Xu and Xie, 2023b; Auer et al., 2023) that only capture temporal dependencies, our approach allows for modeling spatiotemporal dependencies among different time series. Sec. 3.1 presents the details of the proposed conformal inference procedure by assuming that the relational structure at each time step is defined by an adjacency matrix ${\bm{A}}\in{\mathbb{R}}^{N\times N}$ (Sec. 3.1). We then show how to learn the graph structure directly from data and make the model adaptive in Sec. 3.2 and Sec. 3.4, respectively. Finally, we discuss the theoretical properties of the approach in Sec. 3.3. Fig. 1 shows an overview of the architecture.
@@ 3.1 Relational Quantile Predictor
Consider a standard SCP setup, where the training data are split into training and calibration sets. For the moment, we disregard possible nonstationarities in the data considering the problem setup introduced in Sec. 2.1 and encode spatial dependencies in the adjacency matrix ${\bm{A}}\in{\mathbb{R}}^{N\times N}$. While the training set is used to fit the point predictor ${\mathcal{F}}_{\theta}$, we use the prediction residuals in the calibration set (${\mathcal{R}}^{cal}$) to learn the quantile function of the error distribution at each step.
Relational quantile regression
We implement the quantile regressor as a hybrid global-local STGNN, which mixes global (shared) parameters with local, target-specific components (Smyl, 2020). Sharing most learnable parameters across all time series reduces sample complexity, while local parameters allow for tailoring the processing to each series. Specifically, we keep all processing blocks shared and associate a learnable node embedding ${\bm{v}}^{i}\in{\mathbb{R}}^{d_{v}}$ with each time series (Cini et al., 2023b). More specifically, our model is a quantile network (see Sec. 2.3) composed of the following processing layers:
@@
$\hat{\bm{q}}^{i,\alpha}_{t+H}$ $=\textsc{QDec}\left(\alpha,{\bm{z}}_{t}^{i},{\bm{v}}^{i}\right),$ (12)
where ${\bm{r}}^{i}_{t-1}$ are prediction residuals (Eq. 5) and $\hat{\bm{q}}^{\alpha,i}_{t+H}$ is the predicted $\alpha$-quantile at time step $t+H$ for the $i$-th time series. $\textsc{Enc}({}\cdot{})$ denotes any encoding layer, e.g., a linear transformation or an MLP. For the STGNN block, several designs are possible (e.g., see Jin et al. 2023); the one we follow is the template in Sec. 2.4. $\textsc{QDec}({}\cdot{})$ is a readout mapping the representations at each node to the prediction of the quantile of specified level $\alpha$. We refer to the family of quantile networks defined in Eq. 1012 as relational quantile predictors (RelQPs) and use the notation
@@
where ${\mathcal{Q}}_{\psi}$ indicates the shared (global) part of the network and $\widehat{\bm{Q}}^{\alpha}_{t}\in{\mathbb{R}}^{N}$ denotes the predicted $\alpha$-quantiles at time step $t$ w.r.t. the full time series collection. Note that the framework can easily accommodate further inputs at the encoding block (e.g., we can condition the regression on ${\bm{X}}_{<t}$ and ${\bm{U}}_{<t}$). The model is trained by minimizing the pinball loss (Eq. 6) at each time step in the calibration set w.r.t. the full-time series collection. Through the message-passing layers, the residuals of each time series contribute to estimating the quantiles of the error distribution at neighboring nodes. In practice, we restrict the input of the regressor to the most recent observations rather than considering the full sequence (the window length here can also be different from the one used by the point predictor).
@@
$\widehat{C}^{\alpha}_{i,t}=\left[\hat{\bm{x}}^{i}_{t+H}+\hat{{\bm{q}}}^{i,\alpha/2}_{t},\hat{\bm{x}}^{i}_{t+H}+\hat{{\bm{q}}}^{i,1-\alpha/2}_{t}\right],$ (14)
@@
$\hat{\beta}_{i}$ $=\operatorname*{arg\,min}_{\beta_{i}}\Big{|}\hat{{\bm{q}}}^{i,1-\alpha/2+\beta_{i}}_{t}-\hat{{\bm{q}}}^{i,\alpha/2+\beta}_{t}\big{|}$ (15)
$\widehat{C}^{\alpha}_{i,t}$ $=\left[\hat{\bm{x}}^{i}_{t+H}+\hat{{\bm{q}}}^{i,\alpha/2+\hat{\beta}_{i}}_{t},\hat{\bm{x}}^{i}_{t+H}+\widehat{{\bm{q}}}^{i,1-\alpha/2+\hat{\beta}_{i}}_{t}\right],$ (16)
where $\widehat{C}^{\alpha}_{i,t}$ indicates the estimated PI. While both Eq. 14 and Eq. 16 correspond to the same confidence level, the PI in Eq. 16 can be narrower, at the expense of the additional computation needed to obtain $\hat{\beta}$ (Xu and Xie, 2023a). In practice, one can choose between the two approaches given computational constraints and the difference in performance observed on a validation set. Note that $\widehat{C}^{\alpha}_{i,t}$ can provide only approximate coverage, as it is subject to approximation errors of the true quantile function. Sec. 3.3 will discuss this aspect in detail. A potential drawback of CoRel is that residuals cannot be assumed exchangeable in most practical scenarios. The error distribution can be non-stationary, making it difficult to obtain any coverage guarantee. To mitigate the problem, Sec. 3.4 discusses an efficient and scalable approach to make the framework adaptive by updating local components of the architecture over time. Finally, it is worth noting that the use of relational components in CoRel relies on the actual presence of the associated dependencies in the data. In practical applications, the presence of spatiotemporal correlations in the residuals can be verified through ad-hoc statistical tests  (Zambon and Alippi, 2022, 2023), whose outcome can support the adoption of CoRel.